OpenStudy (anonymous):

find f(3) if f(x) =2/5(x+7)

OpenStudy (anonymous):

put the value of x=3 in your given equation of f(X)=2/5(X+7) ie it becomes f(3)=2/5(3+7) =2/5(10) =2/50 =1/25 = 0.04 Therefore,the value of f(3) is 1/25 or 0.04

OpenStudy (anonymous):

ty!!

OpenStudy (anonymous):

wc!!